Calculate the quartiles, (i,e., Q1 and Q3,), deciles D3 and D7, Percentiles P10, P23 and P92 for the


following data. Also find the Quartile Range and Quartile Deviation.


Marks 0-10 10-20 20-30 30-40 40-50 50-60 60-70 70-80


No. of Students 8 7 5 12 28 22 8 10



1
Expert's answer
2022-01-20T12:44:58-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Class & f_i & cf \\
 \hline
 0-10 & 8 & 8 \\ \hdashline
 10-20 & 7 & 15 \\ \hdashline
 20-30 & 5 & 20 \\ \hdashline
 30-40 & 12 & 32 \\ \hdashline
 40-50 & 28 & 60 \\ \hdashline
 50-60 & 22 & 82 \\ \hdashline
 60-70 & 8 & 90 \\ \hdashline
 70-80 & 10 & 100 \\ \hdashline
 & n=100 & \\ 
\end{array}"


1.

"Q_1"  class :

Class with "(\dfrac{n }{4})^{th}" value of the observation  in cf column

"=(\dfrac{100 }{4})^{th}" value of the observation  in cf column

"=(25)^{th}" value of the observation  in cf column

and it lies in the class 30-40.

"Q_1" class : 30-40

The lower boundary point of 30-40 is 30.

"L=30"

"Q_1=L+\dfrac{\dfrac{n }{4}-cf }{f}\cdot c"

"=30+\dfrac{\dfrac{100 }{4}-20 }{12}\cdot 10"

"=34\dfrac{1 }{6}\approx34.1667"


"Q_3"  class :

Class with "(\dfrac{3n }{4})^{th}" value of the observation  in cf column

"=(\dfrac{3\cdot100 }{4})^{th}" value of the observation  in cf column

"=(75)^{th}" value of the observation  in cf column

and it lies in the class 50-60.

"Q_3" class : 50-60

The lower boundary point of 50-60 is 50.

"L=50"

"Q_3=L+\dfrac{\dfrac{3n }{4}-cf }{f}\cdot c"

"=50+\dfrac{\dfrac{300 }{4}-60 }{22}\cdot 10"

"=56\dfrac{9 }{11}\approx56.8182"



2.

"D_3" class :

Class with "(\dfrac{3n }{10})^{th}" value of the observation in cf column

"=(\dfrac{300 }{10})^{th}" value of the observation in cf column

"=(30)^{th}" value of the observation in cf column

and it lies in the class 30-40.

"D_3" class : 30-40

The lower boundary point of 30-40 is 30.

"L=30"

"D_3=L+\dfrac{\dfrac{3n }{10}-cf }{f}\cdot c"

"=30+\dfrac{\dfrac{300 }{10}-20 }{12}\cdot 10"

"=38\dfrac{1 }{3}\approx38.3333"


"D_7" class :

Class with "(\dfrac{7n }{10})^{th}" value of the observation in cf column

"=(\dfrac{700 }{10})^{th}" value of the observation in cf column

"=(70)^{th}" value of the observation in cf column

and it lies in the class 50-60.

"D_7" class : 50-60

The lower boundary point of 50-60 is 50.

"L=50"

"D_7=L+\dfrac{\dfrac{7n }{10}-cf }{f}\cdot c"

"=50+\dfrac{\dfrac{700 }{10}-60 }{22}\cdot 10"

"=54\dfrac{6 }{11}\approx54.5455"


3.

"P_{10}" class :

Class with "(\dfrac{10n }{100})^{th}" value of the observation in cf column

"=(\dfrac{1000 }{100})^{th}" value of the observation in cf column

"=(10)^{th}" value of the observation in cf column

and it lies in the class 10-20.

"P_{10}" class : 10-20

The lower boundary point of 10-20 is 10.

"L=10"

"P_{10}=L+\dfrac{\dfrac{10n }{100}-cf }{f}\cdot c"

"=10+\dfrac{\dfrac{1000 }{100}-8 }{7}\cdot 10"

"=12\dfrac{6}{7}\approx12.8571"


"P_{23}" class :

Class with "(\dfrac{23n }{100})^{th}" value of the observation in cf column

"=(\dfrac{2300 }{100})^{th}" value of the observation in cf column

"=(23)^{th}" value of the observation in cf column

and it lies in the class 30-40.

"P_{23}" class : 30-40

The lower boundary point of 30-40 is 30.

"L=30"

"P_{23}=L+\dfrac{\dfrac{23n }{100}-cf }{f}\cdot c"

"=30+\dfrac{\dfrac{2300 }{100}-20 }{12}\cdot 10"

"=32.5"


"P_{92}" class :

Class with "(\dfrac{92n }{100})^{th}" value of the observation in cf column

"=(\dfrac{9200 }{100})^{th}" value of the observation in cf column

"=(92)^{th}" value of the observation in cf column

and it lies in the class 70-80.

"P_{92}" class : 70-80

The lower boundary point of 70-80 is 70.

"L=70"

"P_{92}=L+\dfrac{\dfrac{92n }{100}-cf }{f}\cdot c"

"=70+\dfrac{\dfrac{9200 }{100}-90 }{10}\cdot 10"

"=72"


3.

Quartile Range "IQR=Q_3-Q_1=56.8182-34.1667=22.6515"


Quartile Deviation"=\dfrac{Q_3-Q_1}{2}=\dfrac{56.8182-34.1667}{2}=11.3258"

Coefficient of Quartile deviation "=\dfrac{Q_3-Q_1}{Q_3+Q_1}"


"=\dfrac{56.8182-34.1667}{56.8182-34.1667}=0.249"